# G2D: From Global to Dense Radiography Representation Learning via Vision-Language Pre-training

### v1

@@ 4.6 Ablation Studies
Pseudo Segmentation vs. Reconstruction  In Tab 2(c), we evaluate the impact of our proposed PS pretext task in comparison to pixel reconstruction and models without a decoder-level constraint. The model pre-trained with PS outperforms the other two approaches across all three downstream tasks, particularly in semantic segmentation. While the model pre-trained with a pixel reconstruction constraint exhibit improved performance compared to unconstrained variants, such models are still outperformed by the model with the PS constraint. These results underscore the effectiveness of decoder-level pretext tasks and suggest that an emphasis on high-level semantics, derived from PS, is more beneficial than focusing on the low-level semantics from pixel reconstruction. The PS potentially reduces the disparity between features learned through VLP and those required by downstream semantic segmentation tasks. It also enables the model to acquire more representative features that are beneficial for various tasks.
Criteria of Pseudo Mask Construction  As shown in Tab 2(d), performance varies with different thresholds, with the 85% quantile threshold proving most effective across all three downstream tasks. Despite employing the Gaussian Mixture Model (GMM) for pseudo mask creation, as suggested by [5], its performance is still surpassed by the 85% quantile approach. This indicates that the original attention map might contain noise, and a higher threshold is beneficial for generating more effective pseudo masks.
Furthermore, Tab 2(f) highlights the importance of aggregating multi-head attention maps for mask construction. Given the absence of explicit semantic supervision in the PS pretext task, not aggregating these maps leads to the creation of multiple pseudo masks. This excess of masks introduce ambiguous/conflicting training objectives for VLP.
Impact of Mask Refinement  Refinement of the pseudo masks affects the model’s efficacy, as shown in Tab 2(h). Performance tends to decrease when either the body mask is omitted or edge smoothing is not applied. However, integrating both these strategies, as we implement in G2D, yields optimal results. This underscores the vital role of pseudo mask refinement in enhancing model performance.
Hyperparameters in VLP  VLP performance is sensitive to hyperparameter choices, particularly the number of attention heads and projector dimensionality. Performance improves with more attention heads, peaking at 3 heads before declining slightly at 4 (Tab 2(g)). Optimal segmentation and classification results occur with 128-dimensional projectors; while 256 dimensions slightly benefit object detection, they lower performance in other tasks (Tab 2(e)). Projectors of 512 dimensions show no further gains. Hence, we select 3 attention heads and 128-dimensional projectors for an optimal balance of complexity and effectiveness.

### v2

@@ 4.6 Ablation Studies
Pseudo Segmentation vs. Reconstruction  In Tab 4(a), we evaluate the impact of our proposed PS pretext task in comparison to pixel reconstruction and models without a decoder-level constraint. The model pre-trained with PS outperforms the other two approaches across all three downstream tasks, particularly in semantic segmentation. While the model pre-trained with a pixel reconstruction constraint exhibit improved performance compared to unconstrained variants, such models are still outperformed by the model with the PS constraint. These results underscore the effectiveness of decoder-level pretext tasks and suggest that an emphasis on high-level semantics, derived from PS, is more beneficial than focusing on the low-level semantics from pixel reconstruction. The PS potentially reduces the disparity between features learned through VLP and those required by downstream semantic segmentation tasks. It also enables the model to acquire more representative features that are beneficial for various tasks.
Criteria of Pseudo Mask Construction  As shown in Tab 4(b), performance varies with different thresholds, with the 85% quantile threshold proving most effective across all three downstream tasks. Despite employing the Gaussian Mixture Model (GMM) for pseudo mask creation, as suggested by [5], its performance is still surpassed by the 85% quantile approach. This indicates that the original attention map might contain noise, and a higher threshold is beneficial for generating more effective pseudo masks.
Furthermore, Tab 4(d) highlights the importance of aggregating multi-head attention maps for mask construction. Given the absence of explicit semantic supervision in the PS pretext task, not aggregating these maps leads to the creation of multiple pseudo masks. This excess of masks introduce ambiguous/conflicting training objectives for VLP.
Impact of Mask Refinement  Refinement of the pseudo masks affects the model’s efficacy, as shown in Tab 4(f). Performance tends to decrease when either the body mask is omitted or edge smoothing is not applied. However, integrating both these strategies, as we implement in G2D, yields optimal results. This underscores the vital role of pseudo mask refinement in enhancing model performance.
Hyperparameters in VLP  VLP performance is sensitive to hyperparameter choices, particularly the number of attention heads and projector dimensionality. Performance improves with more attention heads, peaking at 3 heads before declining slightly at 4 (Tab 4(e)). Optimal segmentation and classification results occur with 128-dimensional projectors; while 256 dimensions slightly benefit object detection, they lower performance in other tasks (Tab 4(c)). Projectors of 512 dimensions show no further gains. Hence, we select 3 attention heads and 128-dimensional projectors for an optimal balance of complexity and effectiveness.